The plans for a rectangular deck call for the width to be 10 feet less than the length. Sam wants the deck to have an overall perimeter of 64 feet. What should the length of the deck be?
Answer by amoresroy(361)

The plans for a rectangular deck call for the width to be 10 feet less than the length. Sam wants the deck to have an overall perimeter of 64 feet. What should the length of the deck be?
Let L = the length of the rectangular deck
W = the width of the rectangular deck
W = L -6
2(L-6) + 2L = 64
2L -12 + 2L = 64
4L = 76
L = 19
The length of the rectangular deck should be 19 feet.